OpenStudy (anonymous):

help? I will fan and medal!! I also want explanation not just answers!!!! Is the simplified form of 2square root of 3 • 2square root of 6 rational? Yes No

OpenStudy (anonymous):

No. \[2\sqrt{3}*2\sqrt{6}=\sqrt{12}*\sqrt{24}=\sqrt{288}\] Because 288 is not a perfect square, this is not rational.

OpenStudy (anonymous):

thank u so much